OpenStudy (anonymous):

What is explicit formula for each term in the sequence such that the pattern is as follows: {5,1,5,1,5,1,5,1...}? No idea how to do this...I tried {5^n/5^(1-n)} obviously doesn't work. I think it is something like this so it differs from odd to even terms...but I can't figure it out.

OpenStudy (zarkon):

\[a_{n}=\alpha+(-1)^{n}\beta\] \[a_0=\alpha+(-1)^0\beta=5\Rightarrow \alpha+\beta=5\] \[a_1=\alpha+(-1)^1\beta=5\Rightarrow \alpha-\beta=1\] add the two equations then \[2\alpha=6\Rightarrow \alpha=3\] then \[3-\beta=1\Rightarrow \beta=2\] so \[a_{n}=3+(-1)^n\cdot 2,\,\,\,n=0,1,2\ldots\]

OpenStudy (anonymous):

how do you know to use a(k) = alpha + (-1)^k * beta?

OpenStudy (zarkon):

because the sequence is just repeating two numbers